Sebio is four times as old as kenneth and sebio is 36yrs older than kenneth.how old is sebio?
Answer by Cromlix(4381)

Hi there,
Sebio is four times as old as kenneth
S = 4K
sebio is 36yrs older than kenneth
S = 36 + K
Substitute 4K for S
4K = 36 + K
Collect like terms:
4K - K = 36
3K = 36
K = 12
Sebio is 48 years old
Kenneth is 12 years old
Hope this helps :-)